Operation Summary for September 20, 2010 - September 22, 2010
Monday Sept 20
The day shift began with Operations monitoring the Tevatron (TeV) store 8099 (60.92E30) and the Recycler stash (372.2E10), with the Antiproton Source (Pbar) stacking (14.0E10), with the Linac, Booster, the and Main Injector (MI) operating normally, with the NuMI off for target work, with MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
During the day shift, Operations terminated store 8099 at 9:03 AM, put the TeV into HEP shot setup, and established store 8100 at 10:07 AM with an average initial luminosity (AIL) of 289.05E30. Operations established beam to Meson MTest experiments T992 and T995 at 10:48 AM. MI personnel accessed MI-31. Operations reported that a TeV separator at sectorC4 sparked and caused a 7E30 loss in luminosity. Operations also reported several Linac RF station trips.
During the evening shift, Operations disabled the MTest beam around 6 PM. Operators halted stacking for about 30 minutes to investigate an AP0 air flow problem; they found that the blower motor had tripped off; it was reset. Operations reported problems with the A1 cryo system; operators reported that the building was very warm. Operators breakered off a heater that was stuck on.
Tuesday Sept. 21
The midnight shift began with Operations monitoring store 8100 (70.75E30) and the stash (341.6E10), with Pbar stacking (6.3E10), and with MiniBooNE taking beam.
During the midnight shift, Operations terminated store 8100 at 1:24 AM and put the TeV into shot setup. During shot setup the TeV quenched at step #24. Operators spent the rest of the shift in quench recovery and investigation. Operations turned off the MiniBooNE beam at 7 AM for an access during the day shift.
During the day shift, TeV personnel tuned and conducted studies. FESS personnel worked on a MiniBooNE MI-12 air conditioning unit: they also accessed the MI-12B enclosure. Operations put the TeV into shot setup and established store 8105 at 2:40 PM with an AIL of 267.28E30. MiniBooNE resumed taking beam.
During the evening shift, Operations reported that the MIRF13 bias power supply tripped its breaker and caused a short downtime; it was reset. Operations disabled the MTest beam at 6:06 PM.
Wednesday Sept. 22
The midnight shift began with Operations monitoring store 8105 (81.8E30) and the stash (232.8E10), with Pbar stacking (23.9E10), and with MiniBooNE taking beam.
During the midnight shift, Operations terminated store 8105 at 2:50 AM, put the TeV into shot setup, and established store 8106 at 4:11 AM with an AIL of 289.1E30. Operations established beam to MTest at 7:48 AM.
The Plans for Wednesday
The plans for today are to stack, stash, and store. The Run Coordinator will allow Linac personnel to access the H- Source dome, a TeV end-of-store study, and a MI study. NuMI personnel reported that their target has a water leak as well as a helium leak; this will delay the NuMI startup by two days.
Comments and Suggestions
|last modified 09/22/2010 email Fermilab|
|Security, Privacy, Legal|